OpenStudy (anonymous):

Solve for x. (4^-2)(4^x+1)(4^3)=(4^5)

OpenStudy (saifoo.khan):

\[\frac{1}{16} \times (4^x+1) = \frac{4^5}{4^3}\]

OpenStudy (saifoo.khan):

\[\frac{1}{16} \times (4^x+1) = 4^2\]

OpenStudy (saifoo.khan):

\[4^x+1 = 16 \times 16\]

OpenStudy (saifoo.khan):

\[4^x = 155.\]

OpenStudy (saifoo.khan):

\[x = \log_{4}155\]\[x = 3.638\]

OpenStudy (saifoo.khan):

I can't believe i did that.

OpenStudy (saifoo.khan):

LOL.

OpenStudy (saifoo.khan):

It looks scary.

OpenStudy (anonymous):

Thank you! I didn't even know how to start it haha